# Self-Supervised Learning for Time Series Analysis: Taxonomy, Progress, and Prospects

### v1

@@ Abstract
Self-supervised learning (SSL) has recently achieved impressive performance on various time series tasks. The most prominent advantage of SSL is that it reduces the dependence on labeled data. Based on the pre-training and fine-tuning strategy, even a small amount of labeled data can achieve high performance. Compared with many published self-supervised surveys on computer vision and natural language processing, a comprehensive survey for time series SSL is still missing. To fill this gap, we review current state-of-the-art SSL methods for time series data in this article. To this end, we first comprehensively review existing surveys related to SSL and time series, and then provide a new taxonomy of existing time series SSL methods. We summarize these methods into three categories: generative-based, contrastive-based, and adversarial-based. All methods can be further divided into ten subcategories. To facilitate the experiments and validation of time series SSL methods, we also summarize datasets commonly used in time series forecasting, classification, anomaly detection, and clustering tasks. Finally, we present the future directions of SSL for time series analysis.
@@ 1 Introduction
With the great success of SSL in CV and NLP, it is appealing to extend SSL to time series data. However, transferring the pretext tasks designed for CV/NLP directly to time series data is non-trivial, and often fails to work in many scenarios. Here we highlight some typical challenges that arise when applying SSL to time series data.. First, time series data exhibit unique properties such as seasonality, trend, and frequency domain information [13, 14, 15]. Since most pretext tasks designed for image or language data do not consider these semantics related to time series data, they cannot be directly adopted. Second, some techniques commonly used in SSL, such as data augmentation, need to be specially designed for time series data. For example, rotation and crop are the commonly used augmentation techniques for image data [16]. However, these two techniques may break the temporal dependency of the series data. Third, most time series data contain multiple dimensions, i.e., multivariate time series. However, useful information usually only exists in a few dimensions, making it difficult to extract useful information in time series using SSL methods for other data types.
@@
This article provides a review of current state-of-the-art SSL methods for time series data. We begin by summarizing recent reviews on SSL and time series data and then propose a new taxonomy from three perspectives: generative-based, contrastive-based, and adversarial-based. The taxonomy is similar to the one proposed by Liu et al. [21] but specifically concentrated on time series data. For generative-based methods, we describe three frameworks: autoregressive-based forecasting, auto-encoder-based reconstruction, and diffusion-based generation. For contrastive-based methods, we divide the existing work into five categories based on how positive and negative samples are generated, including sampling contrast, prediction contrast, augmentation contrast, prototype contrast, and expert knowledge contrast. For each category, we analyze its insights and limitations. Then we sort out and summarize the adversarial-based methods based on two target tasks: time series generation/imputation and auxiliary representation enhancement. The proposed taxonomy is shown in Fig. 1. We conclude this work by discussing possible future directions for time series SSL, including selection and combination of data augmentation, selection of positive and negative samples in SSCL, the inductive bias for time series SSL, theoretical analysis of SSCL, adversarial attacks and robust analysis on time series, time series domain adaption, pretraining and large models for time series, time series SSL in collaborative systems, and benchmark evaluation for time series SSL.
@@
  • New taxonomy and comprehensive review. We provide a new taxonomy and a detailed and up-to-date review of time series SSL. We divide existing methods into ten categories, and for each category, we describe the basic frameworks, mathematical expression, fine-grained classification, and detailed comparison. To the best of our knowledge, this is the first work to comprehensively and systematically review the existing studies of SSL for time series data.
@@
  • Abundant future directions. We point out some key problems in this field from both applicative and methodology perspectives, analyze their causes and possible solutions, and discuss future research directions for time series SSL. We strongly believe that our efforts will ignite further research interests in time series SSL.
The rest of the article is organized as follows. Section 2 provides some review literature on SSL and time series data. Section 3 to Section 5 describe the generation-based, contrastive-based, and adversarial-based methods, respectively. Section 6 lists some commonly used time series data sets from the application perspective. Section 7 discusses several promising directions of time series SSL, and Section 8 concludes the article.
@@ 2 Related Surveys
In this section, several recent reviews on SSL and time series analysis are scrutinized.
@@ 2.1 Surveys on SSL
The surveys on SSL can be categorized by different criteria. In this paper, we outline four widely used criteria: learning paradigms, pretext tasks, components/modules, and downstream tasks. The details are described as follows.
@@ 2.1.1 Learning paradigms
This category focuses on model architectures and training objectives. The SSL methods can be roughly divided into the following categories: generative-based, contrastive-based, and adversarial-based methods. The model architectures of these categories are shown in Fig. 2 and can be summarized as:
Using the learning paradigm as a taxonomy is arguably the most popular among the existing SSL surveys, including [23, 20, 24, 25, 26, 27, 28]. However, not all surveys cover the above three categories. The readers are referred to these surveys for more details. In Table I, we also provide the data modalities involved in each survey, which can help readers quickly find the research work closely related to them.
Figure 2: Learning paradigms of SSL.
2.1.2 Pretext tasks
The pretext task serves as a means to learn informative representations for downstream tasks. Ericsson et al. [29] provides a very comprehensive review of pretext tasks for multiple modalities, including image, video, text, audio, time series, and graph. The various self-supervised pretexts are divided into five broad families: transformation prediction, masked prediction, instance discrimination, clustering, and contrastive instance discrimination. Unlike the learning-paradigm-based criterion, the pretext-task-based criterion is also related to data modality. For example, Jing and Tian [18] summarize the self-supervised feature learning methods on image and video data, and four categories are discussed: generation-based, context-based, free semantic label-based, and cross modal-based, where cross-modal-based methods construct learning task using RGB frame sequence an optical flow sequence, which are unique features in the video.
TABLE I: An overview of recent SSL surveys. The yellow part indicates the proportion of this modality in all six modalities. It can be seen that this article is mainly concerned with SSL on time series data.
@@ 2.1.3 Components and modules
The literature categorizing SSL methods according to their modules and components throughout the pipeline is also an important direction. Jaiswal et al. [17] and Le-Khac et al. [30] sort out the modules and components required in SSL from different perspectives. The pipeline can be divided into four components: positive and negative samples, pretext task, model architecture, and training loss. The basic intuition behind SSCL is to pull positive samples closer and push negative samples away. Therefore, the first step is to construct true positive and negative samples using data transformation or sampling techniques, such as data augmentation and context sampling on time series. It is worth noting that the choice of data augmentation techniques depends on the data modality. For example, the augmentation methods of image data include crop, color distort, resize, and rotate [16], while for time series, the augmentation method involves noise injection, window slicing, and window wrapping [33]. Furthermore, the sample generation method based on context sampling is mostly used for data with temporal dependencies. The second component is pretext tasks, which is a self-supervised task that acts as an important strategy to learn data representations using pseudo-labels [17]. The third component is model architecture, which determines how positive and negative samples are encoded during training. The major categories include end-to-end [16], memory bank [34], momentum encoder [35], and clustering [36].
2.1.4 Downstream tasks
The last way to categorize SSL literature is by downstream tasks. On the topic of speech processing, Latif et al. [31] summarize the development and challenges of self-supervised techniques in three different downstream tasks, namely automatic speech recognition, speaker recognition, and speaker emotion recognition, respectively.
@@ 2.2 Surveys on time series data
The surveys on time series data can be roughly divided into two categories. The first category focuses on different tasks, such as classification [37, 38], forecasting [39, 40, 41, 42], and anomaly detection [43, 44]. These surveys comprehensively sort out the existing methods for each task. The second category focuses on the key components of time series modeling based on deep neural networks, such as data augmentation [33, 45, 46] and model structure [47, 48]. [33] proposed a new taxonomy that divides the existing data augmentation techniques into basic and advanced approaches. [45] also provides a taxonomy and outlines four families: transformation-based methods, pattern mixing, generative models, and decomposition methods. Moreover, both [33] and [45] empirically compare different data augmentation methods for time series classification tasks. [47] systematically reviews transformer schemes for time series modeling from two perspectives: network structure and applications. The taxonomy proposed by Eldele et al. [19] and Deldari et al. [20] is somewhat similar to our proposed taxonomy, i.e., both taxonomies involve time series self-supervised contrastive learning methods. However, our taxonomy provides more detailed categories and more literature in the contrastive-based approach. More importantly, we also thoroughly review a large set of literature for the generative-based and adversarial-based approaches.
@@ 3 Generative-based Methods
In this category, the pretext task is to generate the expected data based on a given view of the data. In the context of time series modeling, the commonly used pretext tasks include using the past series to forecast the future windows or specific time stamps, using the encoder and decoder to reconstruct the input, and forecasting the unseen part of the masked time series. This section sorts out the existing self-supervised representation learning methods in time series modeling from the perspectives of autoregressive-based forecasting, autoencoder-based reconstruction, and diffusion-based generation. It should be noted that autoencoder-based reconstruction task is also viewed as an unsupervised framework. In the context of SSL, we mainly use the reconstruction task as a pretext task, and the final goal is to obtain the representations through autoencoder models.
@@ 3.1 Autoregressive-based forecasting
Given the current time-stamp $t$, the goal of an autoregressive-based forecasting (ARF) task is to forecast a window of length $K$ after time-stamp $t$ using the series before $t$, which can be expressed as:
@@
In the time series modeling with autoregressive-based forecasting task as a pretext task, Recurrent Neural Networks (RNNs) are widely used thanks to their strong capability in spatiotemporal dynamic behavior modeling or sequence prediction [41, 42, 49, 50]. Therefore, it is also naturally applied in the pretext task based on autoregressive forecasting. THOC [51] constructs a self-supervised pretext task for multi-resolution single-step forecasting called Temporal Self-Supervision (TSS). TSS takes the L-layer dilated-RNN with Skip-Connection structure as the model. By setting skip length, it can ensure that the forecasting tasks can be performed with different resolutions at the same time. In addition to RNNs, the forecasting models based on Convolutional Neural Networks (CNNs) also have been developed [52]. Moreover, STraTS [53] first encodes the time series data into triple representations to avoid the limitations of using basic RNN and CNN in modeling irregular and sparse time series data and then builds the transformer-based forecasting model for modeling multivariate medical clinical time series. Graph-based time series forecasting methods can also be used as a self-supervised pretext task. Compared with RNNs and CNNs, Graph Neural Networks (GNNs) can better capture the correlation between variables in multivariate time series data, such as GDN [54].
Unlike the above methods, SSTSC [55] proposes a temporal relation learning prediction task based on the “Past-Anchor-Future” strategy as a self-supervised pretext task. Instead of directly forecasting the values of the future time windows, SSTSC predicts the relationships of the time windows, which can fully mine the temporal relationship structure in the data.
@@ 3.2 Autoencoder-based reconstruction
(a) Autoregressive-based forecasting task
(b) Autoencoder-based reconstruction task
(c) Diffusion-based generation task
@@
The time-step-wise masking randomly selects a certain proportion of time-steps in the series to mask, so the fine-grained information is easier to capture, but it is difficult to learn contextual semantic information in time series. The segment-wise masking randomly selects segments to mask, which allows the model to pay more attention to slow features in the time series, such as trends or high-level semantic information. STEP [71] divided the series into multiple non-overlapping segments of equal length and then randomly selected a certain proportion of the segments for masking. Moreover, STEP pointed out two advantages of using segment-wise masking: the ability to capture semantic information and reduce the input length to the encoder. Different from STEP, Zerveas et al. [72] performed a more complex masking operation on the time series, i.e., the multivariate time series was randomly divided into multiple non-overlapping segments of unequal length on each variable. Variable-wise masking was introduced by Chauhan et al.[73], who defined a new time series forecasting task called variable subset forecast (VSF). In VSF, the time series samples used for training and inference have different dimensions or variables, which may be caused by the absence of some sensor data. This new forecasting task brings the feasibility of self-supervised learning based on variable-wise masking. Unlike the random masking, TARNet [74] considers the pre-trained model based on the random masking strategy irrelevant to the downstream task, which leads to sub-optimal representations. Therefore, TARNet developed a new data-driven masking strategy that uses self-attention score distribution from downstream task training to determine the time steps associated with the downstream task and then masks out these necessary time steps and reconstructs them.
@@
where $\text{KL}(\cdot)$ represents the Kullback-Leibler divergence. The role of the regularization term is to ensure that the learned distribution $P(Z|X)$ is close to the standard normal distribution, thereby regulating the representation of the latent space. The representation learning method based on VAE can model the distribution of each time step to better capture the complex spatiotemporal dependencies and provide better interpretability in time series modeling tasks. For example, InterFusion [77] is a hierarchical VAE that models inter-variable and temporal dependencies in time series data. OmniAnomaly [78] combines VAE and Planar Normalizing Flow to propose an interpretable time series anomaly detection algorithm. In order to better capture the dependencies between different variables in multivariate time series, GRELEN [79] and VGCRN [80] introduce the graph structure and in VAE. In addition to modeling on regular time series, the methods based on VAE have made progress in sparse and irregular time series data representation learning, such as mTANs [81], HetVAE [82] and P-VAE [83]. The latest work attempts to extract seasonal and trend representations in time series data based on VAE. LaST [84] is a disentangled variational inference framework with mutual information constraints. It separates seasonal and trend representations in the latent space to achieve accurate time series forecasting.

### v3

@@ Abstract
Self-supervised learning (SSL) has recently achieved impressive performance on various time series tasks. The most prominent advantage of SSL is that it reduces the dependence on labeled data. Based on the pre-training and fine-tuning strategy, even a small amount of labeled data can achieve high performance. Compared with many published self-supervised surveys on computer vision and natural language processing, a comprehensive survey for time series SSL is still missing. To fill this gap, we review current state-of-the-art SSL methods for time series data in this article. To this end, we first comprehensively review existing surveys related to SSL and time series, and then provide a new taxonomy of existing time series SSL methods by summarizing them from three perspectives: generative-based, contrastive-based, and adversarial-based. These methods are further divided into ten subcategories with detailed reviews and discussions about their key intuitions, main frameworks, advantages and disadvantages. To facilitate the experiments and validation of time series SSL methods, we also summarize datasets commonly used in time series forecasting, classification, anomaly detection, and clustering tasks. Finally, we present the future directions of SSL for time series analysis.
@@ 1 Introduction
With the great success of SSL in CV and NLP, it is appealing to extend SSL to time series data. However, transferring the pretext tasks designed for CV/NLP directly to time series data is non-trivial, and often fails to work in many scenarios. Here we highlight some typical challenges that arise when applying SSL to time series data. First, time series data exhibit unique properties such as seasonality, trend, and frequency domain information [13, 14, 15]. Since most pretext tasks designed for image or language data do not consider these semantics related to time series data, they cannot be directly adopted. Second, some techniques commonly used in SSL, such as data augmentation, need to be specially designed for time series data. For example, rotation and crop are the commonly used augmentation techniques for image data [16]. However, these two techniques may break the temporal dependency of the series data. Third, most time series data contain multiple dimensions, i.e., multivariate time series. However, useful information usually only exists in a few dimensions, making it difficult to extract useful information in time series using SSL methods from other data types.
@@
This article provides a review of current state-of-the-art SSL methods for time series data. We begin by summarizing recent reviews on SSL and time series data and then propose a new taxonomy from three perspectives: generative-based, contrastive-based, and adversarial-based. The taxonomy is similar to the one proposed by Liu et al. [21] but specifically concentrated on time series data. For generative-based methods, we describe three frameworks: autoregressive-based forecasting, auto-encoder-based reconstruction, and diffusion-based generation. For contrastive-based methods, we divide the existing work into five categories based on how positive and negative samples are generated, including sampling contrast, prediction contrast, augmentation contrast, prototype contrast, and expert knowledge contrast. Then we sort out and summarize the adversarial-based methods based on two target tasks: time series generation/imputation and auxiliary representation enhancement. The proposed taxonomy is shown in Fig. 1. We conclude this work by discussing possible future directions for time series SSL, including selection and combination of data augmentation, selection of positive and negative samples in SSCL, the inductive bias for time series SSL, theoretical analysis of SSCL, adversarial attacks and robust analysis on time series, time series domain adaption, pretraining and large models for time series, time series SSL in collaborative systems, and benchmark evaluation for time series SSL.
@@
  • New taxonomy and comprehensive review. We provide a new taxonomy and a detailed and up-to-date review of time series SSL. We divide existing methods into ten categories, and for each category, we describe the basic frameworks, mathematical expression, fine-grained classification, detailed comparison, advantages and disadvantages. To the best of our knowledge, this is the first work to comprehensively and systematically review the existing studies of SSL for time series data.
@@
  • Abundant future directions. We point out key problems in this field from both applicative and methodology perspectives, analyze their causes and possible solutions, and discuss future research directions for time series SSL. We strongly believe that our efforts will ignite further research interests in time series SSL.
The rest of the article is organized as follows. Section 2 provides some review literature on SSL and time series data. Section 3 to Section 5 describe the generation-based, contrastive-based, and adversarial-based methods, respectively. Section 6 lists some commonly used time series data sets from the application perspective. The quantitative performance comparisons and discussions are also provided. Section 7 discusses promising directions of time series SSL, and Section 8 concludes the article.
@@ 2 Related Surveys
In this section, the definition of time series data is first introduced, and then several recent reviews on SSL and time series analysis are scrutinized.
2.1 Definition of time series data
2.1.1 Univariate time series
A univariate time series refers to an ordered sequence of observations or measurements of the same variable indexed by time. It can be defined as $X=(x_{0},x_{1},x_{2},\ldots,x_{t})$, where $x_{i}$ is the point at timestamp $i$. Most often, the measurements are made at regular time intervals.
2.1.2 Multivariate time series
A multivariate time series consists of two or more interrelated variables (or dimensions) that depend on time. It is a combination of multiple univariate time series and can be defined as $\mathbf{X}=\left[X_{0},X_{1},X_{2},\ldots,X_{p}\right]$, where $p$ is the number of variables.
2.1.3 Multiple multivariate time series
Considering the scenario where distinct sets of multivariate time series are concurrently examined. Analyzing such datasets involves studying each set independently and exploring the relationships between different sets. For instance, if we study meteorological data from different cities, each city’s data forms a multivariate time series, collectively resulting in multiple multivariate time series. This can be articulated as $\mathcal{X}=\left\{\mathbf{X}_{0},\mathbf{X}_{1},\mathbf{X}_{2},\ldots,\mathbf{X}_{n}\right\}$, where $n$ is the number of multivariate time series.
@@ 2.2 Surveys on SSL
The surveys on SSL can be categorized by different criteria. In this paper, we outline three widely used criteria: learning paradigms, pretext tasks and components/modules.
@@ 2.2.1 Learning paradigms
This category focuses on model architectures and training objectives. The SSL methods can be roughly divided into the following categories: generative-based, contrastive-based, and adversarial-based methods. The characteristics and descriptions of the above methods can be found in Appendix A. Using the learning paradigm as a taxonomy is arguably the most popular among the existing SSL surveys, including [22, 20, 23, 24, 25, 26, 27]. However, not all surveys cover the above three categories. The readers are referred to these surveys for more details. In Table I, we also provide the data modalities involved in each survey, which can help readers quickly find the research work closely related to them.
TABLE I: An overview of recent SSL surveys on different modalities. The yellow part indicates the proportion of this modality in all six modalities. It can be seen that this article is mainly concerned with SSL on time series data.
@@
2.2.2 Pretext tasks
The pretext task serves as a means to learn informative representations for downstream tasks. Unlike the learning-paradigm-based criterion, the pretext-task-based criterion is also related to data modality. For example, Ericsson et al. [28] provides a very comprehensive review of pretext tasks for multiple modalities, including image, video, text, audio, time series, and graph. The various self-supervised pretexts are divided into five broad families: transformation prediction, masked prediction, instance discrimination, clustering, and contrastive instance discrimination. Jing and Tian [18] summarize the self-supervised feature learning methods on image and video data, and four categories are discussed: generation-based, context-based, free semantic label-based, and cross modal-based, where cross-modal-based methods construct learning task using RGB frame sequence an optical flow sequence, which are unique features in the video. Gui et al. [30] explore four kinds of pretext tasks in computer vision and natural language processing, including context-based methods, contrastive learning methods, generative algorithms, and contrastive generative methods. Essentially, the core of the pretext tasks is how to construct pseudo-supervision signals. Generally speaking, ignoring the differences in data modalities, existing pretext tasks can be roughly summarized into three categories: context prediction, instance discrimination, and instance generation. The main differences and examples are summarized in Table III. It should be noted that here we only list some commonly used pretexts tasks, and some special pretext tasks are not the focus of this article. The details can be found in Appendix B.2.
@@ 2.2.3 Components and modules
The literature categorizing SSCL methods according to their modules and components throughout the pipeline is also an important direction. Jaiswal et al. [17], Le-Khac et al. [29] and Liu et al. [33] sort out the modules and components required in SSL from different perspectives. Specifically, Liu et al. [33] summarizes the research progress of self-supervised contrastive learning on medical time series data. In summary, the pipeline can be divided into four components: positive and negative samples, pretext task, model architecture, and training loss.
The basic intuition behind SSCL is to pull positive samples closer and push negative samples away. Therefore, the first component is to construct positive and negative samples. According to the suggestions of Le-Khac et al. [29], the main methods can be divided into the following categories: multisensory signals, data augmentation, local-global consistency, and temporal consistency. Additional descriptions regarding the characteristics of these categories can be found in Appendix B.1.
The second component is pretext tasks, which is a self-supervised task that acts as an important strategy to learn data representations using pseudo-labels [17]. Pretext tasks have been summarized and categorized in the previous subsection, so repeated content will not be introduced again. The details can be found in Section 2.2.2 and Appendix  B.2.
The third component is model architecture, which determines how positive and negative samples are encoded during training. The major categories include end-to-end [16], memory bank [34], momentum encoder [35], and clustering [36]. More details of these four architectures are summarized in Appendix B.3.
The fourth component is training loss. As summarized in [29], commonly used contrastive loss functions generally include scoring functions (cosine similarity), energy-based margin functions (pair loss and triplet loss), probabilistic NCE-based functions, and mutual information based functions. More details of these loss functions are summarized in Appendix B.4.
@@ 2.3 Surveys on time series data
The surveys on time series data can be roughly divided into two categories. The first category focuses on different tasks, such as classification [37, 38], forecasting [39, 40, 41, 42], and anomaly detection [43, 44]. These surveys comprehensively sort out the existing methods for each task. The second category focuses on the key components of time series modeling based on deep neural networks, such as data augmentation [45, 46, 47, 33], model structure [48, 49, 33]. [45] proposed a new taxonomy that divides the existing data augmentation techniques into basic and advanced approaches. [46] also provides a taxonomy and outlines four families: transformation-based methods, pattern mixing, generative models, and decomposition methods. Moreover, both [45] and [46] empirically compare different data augmentation methods for time series classification tasks. [48] systematically reviews transformer schemes for time series modeling from two perspectives: network structure and applications. Liu et al. [33] provide a comprehensive summary of the various augmentations applied to medical time series data, the architectures of pre-training encoders, the types of fine-tuning classifiers and clusters, and the popular contrastive loss functions. The taxonomies proposed by Eldele et al. [19], Deldari et al. [20] and Liu et al. [33] are somewhat similar to our proposed taxonomy, i.e., three taxonomies involve time series self-supervised contrastive learning methods. However, our taxonomy provides more detailed categories and more literature in the contrastive-based approach. Although the taxonomy proposed by Liu et al. [33] also focuses on time series data. They emphasize discussion of medical time series data, while we focus more on general time series SSL. More importantly, in addition to contrastive-based approaches, we also thoroughly review a large set of literature for the generative-based and adversarial-based approaches.
@@ 3 Generative-based Methods
In this category, the pretext task is to generate the expected data based on a given view of the data. In the context of time series modeling, the commonly used pretext tasks include using the past series to forecast the future windows or specific time stamps, using the encoder and decoder to reconstruct the input, and forecasting the unseen part of the masked time series. This section sorts out the existing self-supervised representation learning methods in time series modeling from the perspectives of autoregressive-based forecasting, autoencoder-based reconstruction, and diffusion-based generation. It should be noted that the autoencoder-based reconstruction task is also viewed as an unsupervised framework. In the context of SSL, we mainly use the reconstruction task as a pretext task, and the final goal is to obtain the representations through autoencoder models. The illustration of the generative-based SSL for time series is shown in Fig. 2. In Appendix C.1 - C.3, the main advantages and disadvantages of three generative-based submethods are summarized. Furthermore, the direct comparison of the three methods is shown in Appendix C.4.
@@ 3.1 Autoregressive-based forecasting
Given the current time step $t$, the goal of an autoregressive-based forecasting (ARF) task is to forecast $K$ future horizons based on $t$ historical time steps, which can be expressed as:
@@
In the time series modeling with autoregressive-based forecasting task as a pretext task, Recurrent neural networks (RNNs) are widely used thanks to their strong capability in spatiotemporal dynamic behavior modeling or sequence prediction [41, 42, 50, 51]. Therefore, it is also naturally applied in the pretext task based on autoregressive forecasting. THOC [52] constructs a self-supervised pretext task for multi-resolution single-step forecasting called Temporal Self-Supervision (TSS). TSS takes the L-layer dilated RNN with skip-connection structure as the model. By setting skip length, it can ensure that the forecasting tasks can be performed with different resolutions at the same time. In addition to RNNs, the forecasting models based on Convolutional neural networks (CNNs) also have been developed [53]. Moreover, STraTS [54] first encodes the time series data into triple representations to avoid the limitations of using basic RNN and CNN in modeling irregular and sparse time series data and then builds the transformer-based forecasting model for modeling multivariate medical clinical time series. Graph-based time series forecasting methods can also be used as a self-supervised pretext task. Compared with RNNs and CNNs, Graph Neural Networks (GNNs) can better capture the correlation among variables and constituent in multivariate time series data, such as GDN [55] and GTS [56]. Graph-augmented normalizing flow (GANF) is another graph-based approach that can model the conditional dependencies among constituent time series [57]. In order to choose a more appropriate model in building time series SSL task, we further give the advantages and disadvantages of these three commonly used models. The details can be found in Appendix D. Unlike the above methods, SSTSC [58] proposes a temporal relation learning prediction task based on the “Past-Anchor-Future” strategy as a self-supervised pretext task. Instead of directly forecasting the values of the future time windows, SSTSC predicts the relationships of the time windows, which can fully mine the temporal relationship in the data.
(a) Autoregressive-based forecasting task
(b) Autoencoder-based reconstruction task
(c) Diffusion-based generation task
@@ 3.2 Autoencoder-based reconstruction
The time-step-wise masking randomly selects a certain proportion of time-steps in the series to mask, so the fine-grained information is easier to capture, but it is difficult to learn contextual semantic information in time series. The segment-wise masking randomly selects segments to mask, which allows the model to pay more attention to slow features in the time series, such as trends or high-level semantic information. STEP [74] divided the series into multiple non-overlapping segments of equal length and then randomly selected a certain proportion of the segments for masking. Moreover, STEP pointed out two advantages of using segment-wise masking: the ability to capture semantic information and reduce the input length to the encoder. Different from STEP, Zerveas et al. [75] performed a more complex masking operation on the time series, i.e., the multivariate time series was randomly divided into multiple non-overlapping segments of unequal length on each variable. Variable-wise masking was introduced by Chauhan et al.[76], who defined a new time series forecasting task called variable subset forecast (VSF). In VSF, the time series samples used for training and inference have different dimensions or variables, which may be caused by the absence of some sensor data. This new forecasting task brings the feasibility of self-supervised learning based on variable-wise masking. Unlike random masking, TARNet [77] considers the pre-trained model based on the masking strategy irrelevant to the downstream task, which leads to sub-optimal representations. TARNet uses self-attention score distribution from downstream task training to determine the time steps that require masking.
@@
where $\text{KL}(\cdot)$ represents the Kullback-Leibler divergence. The role of the regularization term is to ensure that the learned distribution $P(z|x)$ is close to the standard normal distribution, thereby regulating the representation of the latent space. The representation learning method based on VAE can model the distribution of each time step to better capture the complex spatiotemporal dependencies and provide better interpretability in time series modeling tasks. For example, InterFusion [80] is a hierarchical VAE that models inter-variable and temporal dependencies in time series data. OmniAnomaly [81] combines VAE and Planar Normalizing Flow to propose an interpretable time series anomaly detection algorithm. In order to better capture the dependencies between different variables in multivariate time series, GRELEN [82] and VGCRN [83] introduce the graph structure and in VAE. In addition to modeling on regular time series, the methods based on VAE have made progress in sparse and irregular time series data representation learning, such as mTANs [84], P-VAE [85] and HetVAE [86]. The latest work attempts to extract seasonal and trend representations in time series data based on VAE. LaST [87] is a disentangled variational inference framework with mutual information constraints. It separates seasonal and trend representations in the latent space to achieve accurate time series forecasting.